OpenStudy (anonymous):

Will fan and medal! Which statement is true for the equation 3x - 3x - 2 = -2? It has no solution. It has one solution. It has two solutions. It has infinitely many solutions.

OpenStudy (anonymous):

@iGreen

OpenStudy (igreen):

3x - 3x - 2 = -2 Add 2 to both sides, -2 + 2 = ?

OpenStudy (igreen):

@Ezpeasy Please do not give answers, it's against OpenStudy's Code-of-Conduct. We are here to guide the asker to the answer, so we need his/her participation. Please read the \(\bf\color{blue}{\underline{{\href{http:///openstudy.com/code-of-conduct}{Code-of-Conduct}}}}\) to learn more.

OpenStudy (anonymous):

0

OpenStudy (igreen):

Yes, so we have 3x - 3x = 0 What's 3x - 3x?

OpenStudy (anonymous):

x

OpenStudy (igreen):

No, check again.

OpenStudy (igreen):

Anything subtracted by itself equals what?

OpenStudy (anonymous):

0x @iGreen

OpenStudy (anonymous):

@iGreen

OpenStudy (igreen):

It's just 0. That gives us: 0 = 0 So it has infinitely many solutions